OpenStudy (shubhamsrg):

prove that : C(2n,n) / 4^n <= 1/sqrt(3n +1)

OpenStudy (agent0smith):

first part is a combination, right? \[\large C(2n,n) = \frac{ (2n)! }{ n! (2n-n)! }\]

OpenStudy (shubhamsrg):

yes

OpenStudy (shubhamsrg):

@mukushla

OpenStudy (zarkon):

proof by induction works

OpenStudy (shubhamsrg):

oh I tried couldn't make much progress. Please give me any hint ?

OpenStudy (shubhamsrg):

hold on i am trying

OpenStudy (zarkon):

start with the LHS write \[{2(n+1)\choose n+1}\] in terms of \[{2n\choose n}\]

OpenStudy (zarkon):

I'll be back in ~30 min or so..time to walk the dog ;)

OpenStudy (shubhamsrg):

after letting C(2k,k) /4^k <= 1/sqrt(3k +1) to be true. now we check for k+1, in the end we get on LHS (2k+1)/(2k+2) * C(2k,k)/4^k right ?

OpenStudy (shubhamsrg):

hm alright I'll wait

OpenStudy (shubhamsrg):

yes, since (2k+1)/(2k+2) is always <1, hence the statement for k+1 is always true whenever it is true for k thanks a lot sire that was quite easy actually..

OpenStudy (zarkon):

not quite

OpenStudy (zarkon):

\[{2(n+1)\choose n+1}/4^{n+1}=\frac{(2n+2)!}{(n+1)!(2n+2-(n+1))!4^{n+1}}\] \[=\frac{(2n+2)(2n+1)}{(n+1)(n+1)}\frac{1}{4}{2n\choose n}/4^{n}\]

OpenStudy (zarkon):

\[\le\frac{(2n+2)(2n+1)}{(n+1)(n+1)}\frac{1}{4}\frac{1}{\sqrt{3n+1}}=\frac{(n+1)(2n+1)}{(n+1)(n+1)}\frac{1}{2}\frac{1}{\sqrt{3n+1}}\]

OpenStudy (shubhamsrg):

thats what I had said ?

OpenStudy (zarkon):

\[=\frac{(2n+1)}{(n+1)}\frac{1}{2}\frac{1}{\sqrt{3n+1}}\]

OpenStudy (zarkon):

but you need to compare it to \[\frac{1}{\sqrt{3(n+1)+1}}\]

OpenStudy (shubhamsrg):

ooh :P yes o yes

OpenStudy (shubhamsrg):

maybe I got carried away :|

OpenStudy (zarkon):

\[=\frac{(2n+1)}{(n+1)}\frac{1}{2}\frac{1}{\sqrt{3n+1}}\frac{\sqrt{3n+4}}{\sqrt{3n+4}}\]

OpenStudy (zarkon):

\[=\frac{(2n+1)}{(n+1)}\frac{1}{2}\frac{\sqrt{3n+4}}{\sqrt{3n+1}}\frac{1}{\sqrt{3n+4}}\]

OpenStudy (zarkon):

now show \[\frac{(2n+1)}{(n+1)}\frac{1}{2}\frac{\sqrt{3n+4}}{\sqrt{3n+1}}\] is less than 1 for \(n\ge1\)

OpenStudy (shubhamsrg):

why less than one? surely (2n+1)/(2n+2) is less than one but sqrt(3n+4)/sqrt(3n+1) is greater than 1, how did you know there product is less than 1 ?

OpenStudy (zarkon):

because I can show it to be true

OpenStudy (zarkon):

if you show it is less than 1 then you have your proof

OpenStudy (shubhamsrg):

yes. i understand till here

OpenStudy (zarkon):

compute this \[\left[\frac{(2n+1)}{(n+1)}\frac{1}{2}\frac{\sqrt{3n+4}}{\sqrt{3n+1}}\right]^2\]

OpenStudy (zarkon):

then use the fact that if \(a>0\) and \(a^2<1\) then \(a<1\)

OpenStudy (shubhamsrg):

not too sure what difference will it make ? should I open up the brackets ?

OpenStudy (zarkon):

square everything on the inside...

OpenStudy (zarkon):

\[=\frac{(2n+1)^2}{(n+1)^2}\frac{1}{2^2}\frac{{3n+4}}{{3n+1}}=\frac{12n^3+28n^2+19n+4}{12n^3+28n^2+20n+4}\]

OpenStudy (shubhamsrg):

ohh cool.. how did it ever strike you this pattern will come up? :O

OpenStudy (zarkon):

clearly the denominator is bigger

OpenStudy (zarkon):

experience

OpenStudy (shubhamsrg):

well thats great.. surely.. thanks a loads..

OpenStudy (zarkon):

np